Using the "equipartion" of energy concept, estimate the ideal gas $c_p$ (not $c_v$) in units of $k_B$ for "low-temperature" where no vibrational modes are activated. (if you calculate $c_p = 42k_B$, enter 42 as your answer)





1
Expert's answer
2022-03-16T10:19:01-0400

According to "equipartion" of energy concept the gas atom has three degrees of freedom and, therefore, will have an average total energy of 3/2 * Kb * T.

The equation of state for an ideal gas for N particles:

P * V = n * R * T = N * Kb * T.

According to the equation:

C * V * R = N * Kb.

C * V = N * Kb / R = N/NA.